define f : X → Y, g : Y → Z and h : Y → Y
1
Expert's answer
2019-12-02T09:46:18-0500

"f:X \rightarrow Y" defines that f is function which takes in the value from the set of X and return the values from the set of Y.

For example: "f(x)=x^2"

This function takes in the value from the set of all real numbers but returns the value from the set of all positive numbers.

Similarly, "g:Y\rightarrow Z" defines that g is function which takes in the value from the set of Y and return the values from the set of Z.

For example: "g(x)=-\sqrt{x}"

This function takes in the value from the set of all positive numbers but returns the value from the set of all negative numbers.

Similarly, "h:Y\rightarrow Y" defines that h is a function which takes in the value from the set of Y and also returns the values from the set of Y.

For example: "h(x)=\sqrt{x}"

This function takes in the value from the set of all positive numbers and also returns the value from the set of all positive numbers.